OpenStudy (anonymous):

Integral of 0 to (pi/2) of : ((sec^2)(x))/((tan(x)-1)^2) dx Can anyone solve this???

OpenStudy (loser66):

substitute u = tan (x) -1

OpenStudy (anonymous):

What ? Lol

OpenStudy (anonymous):

Oh. Sorry :0

OpenStudy (loser66):

not too late, fix it, friend

OpenStudy (anonymous):

I Did

OpenStudy (anonymous):

hey Loser, here's what I get: integral 1/u^2du = -1/u + C = 1/1-tan(x) + C = sin(x)/cos(x)-sin(x) evaluated at 0 and pi/2 = -1 - 0 = -1 Does this make any sense?

OpenStudy (loser66):

we don't have C, right?

OpenStudy (anonymous):

right, Im just first evaluating as an indefinite integral

OpenStudy (loser66):

I don't know why you have to convert to sin, cos

OpenStudy (anonymous):

just thought it would be easier to solve, same answer in either case?

OpenStudy (anonymous):

is the answer even plausible?

OpenStudy (loser66):

oh yea, you have to, sorry friend,

OpenStudy (loser66):

so it turns \(\dfrac{cosx}{1-sinx}\)from 0 to \(\tfrac{\pi}{2}\)right?

OpenStudy (anonymous):

actually, cos(x)/(cos(x)-sin(x))

OpenStudy (loser66):

\[\frac {cos \tfrac{\pi}{2}}{1-sin\frac{\pi}{2}}- \frac{cos0}{1-sin0}\]

OpenStudy (loser66):

=-1

OpenStudy (anonymous):

but is that a real answer?

OpenStudy (anonymous):

there's an asymptot in the middle if we graph the function

OpenStudy (loser66):

hehe... you got it before me, what do you mean by "real answer"? is there any fake answer?

OpenStudy (loser66):

what is the original function? you just ask for taking integral, how can I know what happen to the original one? Even I know the original function, I may NOT know what happen with it.